# Robustly optimal dynamics for active matter reservoir computing

### v2

@@ III.8 The optimal regime is robust against variations in homing force strength
Figure 9: Viscoelastic droplets forming in the optimal regime are remarkably robust against variations in homing force strength, under driven conditions. (a) Predictive performance for varied homing force strength $K_{h}\in[10^{-3};10^{3}]$ (global attraction to the center of the simulation box) and speed controller strength $K_{sc}\in[10^{-5};10^{2}]$, using a fixed target agent speed $s=0.04833$. This target agent speed corresponds to Fig. 2, triangle symbol. The optimal homing strength regime in the critically damped regime (triangle symbol) yields viscoelastic soft matter droplets and spans about two orders of magnitude ($1.0\lessapprox K_{h}\lessapprox 20.0$). (b) Snapshots corresponding to parameter combinations marked as symbols in sub-figure (a): underdamped (circle), highly condensed (square), viscoelastic droplet (triangle), loosely bound crystal under near-zero homing force conditions (diamond). Refer to Tab. 13 for corresponding videos.
@@
In Fig. 9a we vary the homing force strength $K_{h}$ for a fixed target agent speed $s$ of the speed-controller, and a varied speed-controller strength $K_{sc}$. The optimal performance $P=0.889$ is reached at $(K_{sc},K_{h})=(0.02069,6.158)$ (triangle symbol). This performance is almost identical to those with a homing force strength $K_{h}=2.0$, presented in Fig. 2a (triangle symbol). The corresponding snapshot in Fig. 9b and the Supplementary Video Tab. 13/2 reveal that the soft matter system forms the previously studied condensed, viscoelastic droplet with an interface enclosing the driver. We note that high performances $P>0.85$ are stable within about two orders of magnitude of varying the homing force strength, underlining its robustness. Lowering the speed-controller strength (circle symbol) recovers the underdamped soft matter regime with a predictive performance $P<0.60$ (see also Fig. 2, cross symbol). Increasing the homing force strength in the critically damped regime (square symbol) leads to dominating homing forces, which cause a strong compression of the soft matter system. Here, agents barely react to the repulsive driver force, resulting in a decrease in performance. Despite this weak driver influence, a prediction score of $P\approx 0.60$ can still be achieved.
@@ Appendix A Chaotic Time-Series Benchmarking Details
In this paper, we use some common chaotic attractors as benchmark systems for the time series prediction task of our active matter reservoir computer (see Sect. II.2). In our default implementation, we integrate the coupled differential equations of the Lorenz-63 system [88] ($\dot{x}=a(y-x);\dot{y}=x(b-z)-y;\dot{z}=xy-cz$) with standard parameters ($a=10;b=28;c=8/3$) using Euler integration with an integration time step $\Delta t=0.02$ and initial conditions $(x_{0},y_{0},z_{0})=(0.0,1.0,1.05)$ for training and $(0.0,1.1,-1.5)$ for testing. For the other chaotic attractors (Hénon-Heiles, Rössler, Chua, Lorenz-96), we generate trajectories using the dysts library (version 0.6) [86] (corresponding publications: Refs. [85, 31]). To compare predictive performance, we choose to normalize these time series to the mean effective driver speed
@@
[…]
@@
[…]
@@ D.3 Speed-controller Parameter Scan
Figure 16: Predictive performance for a speed-controller parameter scan with a smaller integration time step of $\Delta t=0.002$ instead of the default integration time step $\Delta t=0.02$ used in Fig. 2a. Compared to Fig. 2a, there is no sharp drop in predictive performance at the lower dashed yellow line. The arrested regime is shifted outwards towards the lower right-hand side of the diagram, expanding the range for predictive performances $>0.75$ for the Lorenz-63 prediction. Here, the superposition of oscillations (alternating forward-backward motion) observed for the original timestep of $\Delta t=0.02$, disappears for the circle and nabla symbols. We also observe a higher performance beyond the upper yellow dashed line in the upper left-hand side of the diagram. This could be the effect of reduced variations within the Lorenz-63 trajectory integrated with the smaller integration time step, which yields a slightly simpler prediction task.
@@ D.4 Intrinsic Agent Relaxation Dynamics
Figure 20: Intrinsic agent relaxation dynamics for simulations without (a,b) and with (c-f) agent-agent interactions in place and with different integration time steps analogously to Fig. 3a,b. The smaller integration time step of $\Delta t=0.002$ allows resolving relaxation dynamics in the far overdamped regime (nabla symbol) that is otherwise rendered as an arrested regime for $\Delta t=0.02$. The presence of agent-agent interactions (d,f) (which also applies to (c,e)) prominently hinders the oscillations in the relaxation dynamics above critical damping (square, diamond, cross), and subtly shifts the time-points of relaxation for the near-critical and overdamped regimes (triangle, circle).
@@ D.5 Speed-controller Parameter Scans for Different Chaotic Time Series Benchmarks
Figure 24: Dynamical susceptibility for various chaotic time-series benchmarks presented in Fig. 6 and analogous to Fig. 5a for the Lorenz-63 attractor in the main text. High dynamical susceptibility correlates with the near-critically damped high-performance region in Fig. 6 for swarms with low rotation, which is displayed in Fig. 22. Note that rotations and dilatations were not taken into account when computing the velocity fluctuations as a prerequisite for the dynamical susceptibility (see Eq. 16 and its discussion). Because of this, we measure high dynamical susceptibilities $\chi>15.0$ in the underdamped regimes (upper left-hand triangle) of the Hénon–Heiles and Rössler attractors, which feature high rotation as shown in Fig. 22a,b.
@@ Appendix E Supplementary Videos
2 [Uncaptioned image] Speed controller scan (triangle symbol)
@@
2 [Uncaptioned image] Speed controller scan (triangle symbol)
@@
2 [Uncaptioned image] Damping analysis, non-interacting (triangle symbol)
@@
2 [Uncaptioned image] Damping analysis, interacting (triangle symbol)
@@
1 [Uncaptioned image] Alignment force scan (triangle symbol)
@@
2 [Uncaptioned image] Homing force scan (triangle symbol)

### v3

@@ III.8 The optimal regime is robust against variations in homing force strength
Figure 9: Viscoelastic droplets forming in the optimal regime are remarkably robust against variations in homing force strength, under driven conditions. (a) Predictive performance for varied homing force strength $K_{h}\in[10^{-3};10^{3}]$ (global attraction to the center of the simulation box) and speed controller strength $K_{sc}\in[10^{-5};10^{2}]$, using a fixed target agent speed $s=0.04833$. This target agent speed corresponds to Fig. 2, pyramid symbol. The optimal homing strength regime in the critically damped regime (pyramid symbol) yields viscoelastic soft matter droplets and spans about two orders of magnitude ($1.0\lessapprox K_{h}\lessapprox 20.0$). (b) Snapshots corresponding to parameter combinations marked as symbols in sub-figure (a): underdamped (circle), highly condensed (square), viscoelastic droplet (pyramid), loosely bound crystal under near-zero homing force conditions (diamond). Refer to Tab. 13 for corresponding videos.
@@
In Fig. 9a we vary the homing force strength $K_{h}$ for a fixed target agent speed $s$ of the speed-controller, and a varied speed-controller strength $K_{sc}$. The optimal performance $P=0.889$ is reached at $(K_{sc},K_{h})=(0.02069,6.158)$ (pyramid symbol). This performance is almost identical to those with a homing force strength $K_{h}=2.0$, presented in Fig. 2a (pyramid symbol). The corresponding snapshot in Fig. 9b and the Supplementary Video Tab. 13/2 reveal that the soft matter system forms the previously studied condensed, viscoelastic droplet with an interface enclosing the driver. We note that high performances $P>0.85$ are stable within about two orders of magnitude of varying the homing force strength, underlining its robustness. Lowering the speed-controller strength (circle symbol) recovers the underdamped soft matter regime with a predictive performance $P<0.60$ (see also Fig. 2, cross symbol). Increasing the homing force strength in the critically damped regime (square symbol) leads to dominating homing forces, which cause a strong compression of the soft matter system. Here, agents barely react to the repulsive driver force, resulting in a decrease in performance. Despite this weak driver influence, a prediction score of $P\approx 0.60$ can still be achieved.
@@ Appendix A Chaotic Time-Series Benchmarking Details
In this paper, we use some common chaotic attractors as benchmark systems for the time series prediction task of our active matter reservoir computer (see Sect. II.2). In our default implementation, we integrate the coupled differential equations of the Lorenz-63 system [98] ($\dot{x}=a(y-x);\dot{y}=x(b-z)-y;\dot{z}=xy-cz$) with standard parameters ($a=10;b=28;c=8/3$) using Euler integration with an integration time step $\Delta t=0.02$ and initial conditions $(x_{0},y_{0},z_{0})=(0.0,1.0,1.05)$ for training and (0.0,1.1,1.5)(0.0,1.1,-1.5) for testing. For the other chaotic attractors (Hénon-Heiles, Rössler, Chua, Lorenz-96), we generate trajectories using the dysts library (version 0.6) [96] (corresponding publications: Refs. [95, 29]). To compare predictive performance, we choose to normalize these time series to the mean effective driver speed
@@
Then each dysts trajectory is generated using the following procedure: We first generate a fine-grained trajectory with 105 points per period and 2 $\cdot$ 108 steps in total. The trajectory is then fitted to a square box (the driver box) of size 8.0 and centered in a square box (the simulation box) of size 16.0. This ensures sufficient driver interaction with typical swarm sizes of $N=200$ particles. We then determine the best sampling rate to sample the fine-grained trajectory using a simple bisection algorithm. This yields a driver trajectory with the above-determined desired mean speed $\bar{v}_{d}$, with a precision of 10-2. The final trajectories correspond to the sampled fine-grained trajectories with the following allocation: 103 (burn-in steps training) + 5 $\cdot$ 104 (simulation steps training) + 103 (burn-in steps testing) + 5 $\cdot$ 104 (simulation steps testing) steps. We plot the first 103 simulation data points (excluding the burn-in period) of these systems as insets in Fig. 6 and provide videos of the raw trajectories in Tab. 10. We confirm that for our test trajectories of 5 $\cdot$ 104 data points each, the mean driver speeds for the chosen subset of the total time series generated are reasonably close to those of the (non-dysts-generated) Lorenz-63 trajectory: 10.20 $\pm$ 2.78 (Hénon–Heiles), 10.30 $\pm$ 3.84 (Rössler), 9.91 $\pm$ 4.57 (Chua), and 10.31 $\pm$ 6.05 (Lorenz-96). We quantify the chaoticity of the (non-rescaled) attractors using the highest estimated Lyapunov exponents (maximum_lyapunov_estimated) recorded in the provided database chaotic_attractors.json of the dysts package (version 0.6) [96]: $\lambda_{max}^{\text{Hénon–Heiles}}=0.03551120315239528$, $\lambda_{max}^{\text{Rössler}}=0.15059688939547888$, $\lambda_{max}^{\text{Chua}}=1.2301163591725595$, and $\lambda_{max}^{\text{Lorenz-96}}=1.3361667787286362$.
@@
$\Delta t_{\text{pred}}$ $\approx 0.50\ (\text{L63})$
$\approx 12.76\ (\text{Hénon–Heiles})$ $\approx 3.00\ (\text{Rössler})$
$\approx 0.36\ (\text{Chua})$ $\approx 0.34\ (\text{L96})\,.$
@@ D.3 Speed-controller Parameter Scan
Figure 16: Predictive performance for a speed-controller parameter scan with a smaller integration time step of $\Delta t=0.002$ instead of the default integration time step $\Delta t=0.02$ used in Fig. 2a. Compared to Fig. 2a, there is no sharp drop in predictive performance at the lower dashed yellow line. The arrested regime is shifted outwards towards the lower right-hand side of the diagram, expanding the range for predictive performances >0.75>0.75 for the Lorenz-63 prediction. Here, the superposition of oscillations (alternating forward-backward motion) observed for the original timestep of $\Delta t=0.02$, disappears for the circle and nabla symbols. We also observe a higher performance beyond the upper yellow dashed line in the upper left-hand side of the diagram. This could be the effect of reduced variations within the Lorenz-63 trajectory integrated with the smaller integration time step, which yields a slightly simpler prediction task.
@@ D.4 Intrinsic Agent Relaxation Dynamics
Figure 20: Intrinsic agent relaxation dynamics for simulations without (a,b) and with (c-f) agent-agent interactions in place and with different integration time steps analogously to Fig. 3a,b. The smaller integration time step of $\Delta t=0.002$ allows resolving relaxation dynamics in the far overdamped regime (nabla symbol) that is otherwise rendered as an arrested regime for $\Delta t=0.02$. The presence of agent-agent interactions (d,f) (which also applies to (c,e)) prominently hinders the oscillations in the relaxation dynamics above critical damping (square, diamond, cross), and subtly shifts the time-points of relaxation for the near-critical and overdamped regimes (pyramid, circle).
@@ D.5 Speed-controller Parameter Scans for Different Chaotic Time Series Benchmarks
Figure 24: Dynamical susceptibility for various chaotic time-series benchmarks presented in Fig. 6 and analogous to Fig. 5a for the Lorenz-63 attractor in the main text. High dynamical susceptibility correlates with the near-critically damped high-performance region in Fig. 6 for swarms with low rotation, which is displayed in Fig. 22. Note that rotations and dilatations were not taken into account when computing the velocity fluctuations as a prerequisite for the dynamical susceptibility (see Eq. 16 and its discussion). Because of this, we measure high dynamical susceptibilities $\chi>15.0$ in the underdamped regimes (upper left-hand pyramid) of the Hénon–Heiles and Rössler attractors, which feature high rotation as shown in Fig. 22a,b.
@@ Appendix E Supplementary Videos
2 [Uncaptioned image] Speed controller scan (pyramid symbol)
@@
2 [Uncaptioned image] Speed controller scan (pyramid symbol)
@@
2 [Uncaptioned image] Damping analysis, non-interacting (pyramid symbol)
@@
2 [Uncaptioned image] Damping analysis, interacting (pyramid symbol)
@@
1 [Uncaptioned image] Alignment force scan (pyramid symbol)
@@
2 [Uncaptioned image] Homing force scan (pyramid symbol)